# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ Abstract
In the realm of time series analysis, tackling the phenomenon of concept drift poses a significant challenge. Concept drift – characterized by the evolving statistical properties of time series data, affects the reliability and accuracy of conventional analysis models. This is particularly evident in co-evolving scenarios where interactions among variables are crucial. This paper presents Drift2Matrix, a novel framework that leverages kernel-induced self-representation for adaptive responses to concept drift in time series. Drift2Matrix employs a kernel-based learning mechanism to generate a representation matrix, encapsulating the inherent dynamics of co-evolving time series. This matrix serves as a key tool for identification and adaptation to concept drift by observing its temporal variations. Furthermore, Drift2Matrix effectively identifies prevailing patterns and offers insights into emerging trends through pattern evolution analysis. Our empirical evaluation of Drift2Matrix across various datasets demonstrates its effectiveness in handling the complexities of concept drift. This approach introduces a novel perspective in the theoretical domain of co-evolving time series analysis, enhancing adaptability and accuracy in the face of dynamic data environments.
@@ 1 Introduction
Co-evolving time series data analysis plays a crucial role in diverse sectors including finance, healthcare, and meteorology. Within these areas, multiple time series evolve simultaneously and interact with one another, forming complex, dynamic systems. The evolving statistical properties of such data present significant analytical challenges. A particularly pervasive issue is concept drift Lu et al. (2018b); Yu et al. (2024), which refers to shifts in the underlying data distribution over time, thereby undermining the effectiveness of static models. Miyaguchi & Kajino (2019); You et al. (2021).
Traditional time series approaches commonly rely on the assumptions of stationarity and linear relationships. Methods such as ARIMA and VAR Box (2013), for instance, perform well in circumstances with stable and predictable dynamics. However, their effectiveness decreases when dealing with non-stationary data, particularly in the presence of concept drift. Conversely, machine learning methodologies Li et al. (2022); Wen et al. (2020), such as diverse neural network architectures Ho et al. (2022); Li et al. (2023); Yang et al. (2024), offer more flexibility but often require large amounts of data and face difficulties in terms of interpretability and adaptability, especially in dynamic contexts.
@@
The evolving study has steered the field towards more adaptive and dynamic models. Methods like change point detection Deldari et al. (2021); Liu et al. (2023) and online learning algorithms Huang et al. (2022); Zhang et al. (2024) are designed to detect shifts in patterns. Nonetheless, these methods are typically restricted to detecting structural breaks or focusing on univariate series, rather than tracking and predicting subtle, ongoing changes in concepts, which limits their applicability in real-world co-evolving time series. In the complex environments, where multiple time series evolve and interact simultaneously, capturing the nonlinear relationships among variables is criticalMarcotte et al. (2023); Bayram et al. (2022). Despite recent advancements Matsubara & Sakurai (2019); Li et al. (2022); Wen et al. (2024), most multivariate models define the concept as a collective behavior of streaming data, falling short in their ability to capture the dynamics of individual series and their interactions. This limitation impairs the models’ capability to discover and interpret the complex interdependencies among variables, thereby constraining their effectiveness in scenarios that involve multiple time series, such as city-wide electricity usage or road traffic forecasting. This raises a critical question – Can we identify underlying concepts from co-evolving time series and leverage their nonlinear relationships to predict concepts that have not appeared in a single series?
@@
To tackle these challenges, we introduce Drift2Matrix, a framework designed for the dynamic complexities of co-evolving time series data. Drift2Matrix employs a kernel-induced self-representation method, adept at capturing the intricate interdependencies and the evolving natures of such data. Our approach, which transforms time series into a matrix format capable of adapting to concept drift, offers a robust and flexible solution for co-evolving time series analysis amidst non-linear interactions and shifting distributions.
@@
(O1) Concept identification: Fig. 1 (b) displays the heatmap of Drift2Matrix-generated representation matrix ($n\times n$) over time [^1] Drift2Matrix autonomously determines the optimal segments in a domain-agnostic manner.. This matrix showcases a block diagonal structure, where each block represents a unique concept (e.g., C1 - C5), with brightness in the heatmap indicating series correlation strength. Variations of the block structure over time underscore the dynamic correlations within the time series, indicating the presence and evolution of concepts. An extended analysis of these matrices facilitates the identification of the total number of concepts and their pattern (Fig. 1 (c)).
(O2) Dynamic concept drift: Tracking the transitions of time series within the Drift2Matrix-generated matrices provides insights into the trajectory of concept drift. Fig. 1 (d) illustrates the concept drift process over time for each of the 3 example series. Red dashed lines in the figure mark the points where concept drift occurs, exemplified by the transition of Series 1 from Concept 4 to Concept 2 and subsequent shifts. This process can also be observed in the representation matrix shown in Fig. 1 (b), where the red star marks the position of Series 1 (S1), and the purple dashed lines trace how S1 shifts over time, further illustrating the process of tracking its transition between concepts.
(O3) Forecasting: The grey areas in Fig. 1 (d) represent the forecasted series distribution and values. These forecasts, denoted by dashed lines matching the colors of the concepts, align closely with the actual series trend (solid grey lines). A significant forecast for Series 1 includes the emergence of Concept 1, previously undetected. This predictive capability stems from Drift2Matrix’s ecosystem perspective of the co-evolving time series, leveraging a probabilistic model of the nonlinear interactions among series to anticipate the emergence and evolution of new concepts.
Contributions. Drift2Matrix represents more than a mere solution – it is a paradigm shift in co-evolving time series analysis. This framework introduces a new perspective on modeling and interpreting complex, interrelated co-evolving time series. Drift2Matrix has the following desirable properties:
1. Drift2Matrix introduces a novel, kernel-induced approach for modeling complex interdependencies, enhancing understanding of underlying dynamics, and can be easily integrated into most deep learning backbones.
2. Drift2Matrix is adaptive, with the capability to identify and respond to concept drift autonomously, without prior knowledge about concept.
3. Drift2Matrix transforms concept drift into a matrix optimization problem and enhances interpretability, offering a new perspective into the evolving dynamics of co-evolving time series.
@@ 2 The landscape of Concept Drift
Concept Drift. Concept drift in time series refers to the scenario where the statistical properties of the target variable, or the joint distribution of the input-output pairs, change over time. These drifts primarily exhibit in two manners Ren et al. (2018); Kim et al. (2021): the first is characterized by subtle, ongoing changes, reflecting the evolving dynamics of the time series, while the second arises from sudden shifts caused by structural breaks in the relationships among time series. Both gradual and abrupt changes can significantly disrupt model performance if not detected and adapted to in a timely manner, as they challenge the stability and accuracy of predictive models.
Challenges in Co-evolving Scenarios. Recent advances in time series analysis have led to progress in addressing concept drift. Dish-TS Fan et al. (2023) offers a general approach for alleviating distribution shift in time series forecasting by normalizing model inputs and outputs to better handle distribution changes. Similarly, Cogra’s application of the Sequential Mean Tracker (SMT) adjusts to changes in data distribution, improving forecast accuracy Miyaguchi & Kajino (2019). Despite these strides, these methodologies exhibit limitations when applied to co-evolving time series, where interdependencies between series introduce additional complexity. In such scenarios, a shift in one variable can propagate through the network of interrelations, affecting the entire system. DDG-DA Li et al. (2022) for data distribution generation has been adapted to better suit co-evolving scenarios, addressing the unique challenges presented by the interplay of multiple data streams under concept drift conditions. However, this method defines the concept as a collective behavior represented by co-evolving time series rather than capturing the dynamics of individual series and their interactions. Notably, even the most recent deep learning methods that mention concept drift, such as OneNet Wen et al. (2024) and FSNet Pham et al. (2022), primarily aim to mitigate the impact of concept drift on forecasting rather than addressing the challenges of adaptive concept identification and dynamic concept drift. They achieve this by incorporating an ensemble of models with diverse data biases or by refining network parameters for better adaptability. Due to the space limit, more related works about concept-drift, representation learning on times series and motivation are left in the Appendix A.
@@ 3 Preliminaries
Concept. Throughout this paper, a concept is defined as the profile pattern of a cluster of similar subseries, observed within a specific segement/window. Here, the term “profile pattern” refers to a subseries, the vector representation of which aligns with the centroid of similar subseries. We use a tunable hyp erparameter $\rho$, to differentiate profile patterns and modulate whether concept drift is gradual (smaller $\rho$, more concepts) or abrupt (larger $\rho$, fewer concepts). For a more detailed mathematical definition, please refer to Appendix A.
Notation. We denote matrices by boldface capital letters, e.g., $\mathbf{M}$. $\mathbf{M}^{\mathrm{T}}$, $\mathbf{M}^{-1}$, Tr$(\mathbf{M})$ indicate the transpose, inverse and trace of matrix $\mathbf{M}$, respectively. $\mathrm{diag}(\mathbf{M})$ refers to a vector with its $i$-$th$ element being the $i$-$th$ diagonal element of $\mathbf{M}$.
@@ 3.1 Self-representation Learning in Time Series
Time series often manifest recurring patterns. One feasible way to capture these inherent patterns is through self-representation learning Bai & Liang (2020). This approach models each series as a linear combination of others, formulated as $\mathbf{S=SZ}$ or $S_{i}=\sum_{j}S_{j}Z_{ij}$, where $\mathbf{Z}$ is the self-representation coefficient matrix. In multiple time series, high $Z_{ij}$ values indicate similar behaviors or concepts between $S_{i}$ and $S_{j}$. The learning objective function is:
@@ 3.2 Kernel Trick for Modeling Time Series
Addressing nonlinear relationships in co-evolving time series, especially in the presence of concept drift, can be challenging for linear models. Kernelization techniques overcome this by mapping data into higher-dimensional spaces using suitable kernel functions. This facilitates the identification of concepts within these transformed spaces. The process is facilitated by the “kernel trick”, which employs a nonlinear feature mapping, $\Phi(\mathbf{S})$: $\mathcal{R}^{d}\rightarrow\mathcal{H}$, to project data $\mathbf{S}$ into a kernel Hilbert space $\mathcal{H}$. Direct knowledge of the transformation $\Phi$ is not required; instead, a kernel Gram matrix $\bm{\mathcal{K}}=\Phi(\mathbf{S})^{\top}\Phi(\mathbf{S})$ is used.
4 Drift2Matrix
This section introduces the fundamental concepts and design philosophy of Drift2Matrix. Our objective is to identify significant concept trends and encapsulate them into a succinct yet powerful and adaptive representative model.
@@ 4.1 Kernel-Induced Representation Learning
$\min_{\mathbf{Z}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{Z}||^{2}=\min_{\mathbf{Z}}\frac{1}{2}\mathrm{Tr}(\bm{\mathcal{K}}-\alpha\bm{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\bm{\mathcal{K}}\mathbf{Z}),\quad s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$ (2)
Here, the mapping function $\mathrm{\Phi}(\cdot)$ needs not be explicitly identified and is typically replaced by a kernel $\bm{\mathcal{K}}$ subject to $\bm{\mathcal{K}}=\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})$. It’s noteworthy that the parameter $\alpha$ is key to preserving the local manifold structure of time series during this projection, further explained in Sec. 5.2.
@@
$\min_{\mathbf{Z}}\frac{1}{2}\mathrm{Tr}(\bm{\mathcal{K}}-$ $\alpha\bm{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\bm{\mathcal{K}}\mathbf{Z})+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},\quad s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$ (3)
where $\gamma>0$ balances the loss function with regularization term, $||\mathbf{Z}||_{\boxed{\scriptstyle k}}=\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ and $\lambda_{i}(\mathbf{L_{Z}})$ contains the eigenvalues of Laplacian matrix $\mathbf{L_{Z}}$ corresponding to $\mathbf{Z}$ in decreasing order. Here, the regularization term is equal to 0 if and only if $\mathbf{Z}$ is $k$-block diagonal (see Theorem 4.1 for details). Based on the learned high-quality matrix $\mathbf{Z}$ (containing the block diagonal structure), we can easily group the time series into $k$ concepts using traditional spectral clustering technology Ng et al. (2001). The detailed method of estimating the number of concepts $k$ is provided in Appendix B. To solve Eq. 3, which is a nonconvex optimization problem, we leverage the Augmented Lagrange method with Alternating Direction Minimization strategy to propose a specialized method for solving the nonconvex kernel self-representation optimization (see Appendix C.3).
Theorem 4.1
@@
Proof 4.2
Due to the fact that $\mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0$, the corresponding Laplacian matrix $\mathbf{L_{Z}}$ is positive semidefinite, i.e., $\mathbf{L_{Z}}\succeq 0$, and thus $\lambda_{i}(\mathbf{L_{Z}})\geq 0$ for all $i$. The optimal solution of $\min\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ is that all elements of $\lambda_{i}(\mathbf{L_{Z}})$ are equal to 0, which means that the $k$ smallest eigenvalues are 0. Combined with the Laplacian matrix property, it’s evident that the multiplicity $k$ of the zero eigenvalues in Laplacian matrix $\mathbf{L_{Z}}$ matches the count of connected components (or blocks) present in $\mathbf{Z}$, and thus the soundness of Theorem 1 has been proved.

### v3

# Kunpeng Xu    Lifei Chen    Shengrui Wang

@@ Abstract
In the realm of time series analysis, tackling the phenomenon of concept drift poses a significant challenge. Concept drift – characterized by the evolving statistical properties of time series data, affects the reliability and accuracy of conventional analysis models. This is particularly evident in co-evolving scenarios where interactions among variables are crucial. This paper presents CORAL, a simple yet effective method that models time series as an evolving ecosystem to learn representations of concept drift. CORAL employs a kernel-induced self-representation learning to generate a representation matrix, encapsulating the inherent dynamics of co-evolving time series. This matrix serves as a key tool for identification and adaptation to concept drift by observing its temporal variations. Furthermore, CORAL effectively identifies prevailing patterns and offers insights into emerging trends through pattern evolution analysis. Our empirical evaluation of CORAL across various datasets demonstrates its effectiveness in handling the complexities of concept drift. This approach introduces a novel perspective in the theoretical domain of co-evolving time series analysis, enhancing adaptability and accuracy in the face of dynamic data environments, and can be easily integrated into most deep learning backbones.
Machine Learning, ICML
@@ 1 Introduction
Co-evolving time series data analysis plays a crucial role in diverse sectors including finance, healthcare, and meteorology. Within these areas, multiple time series evolve simultaneously and interact with one another, forming complex, dynamic systems. The evolving statistical properties of such data present significant analytical challenges. A particularly pervasive issue is concept drift (Lu et al., 2018b; Yu et al., 2024), which refers to shifts in the underlying data distribution over time, thereby undermining the effectiveness of static models (You et al., 2021).
Traditional time series approaches commonly rely on the assumptions of stationarity and linear relationships. Methods such as ARIMA and VAR (Box, 2013), for instance, perform well in circumstances with stable and predictable dynamics but struggle with non-stationary data, particularly in the presence of concept drift. Deep learning models have made significant strides across various time series tasks. However, a key challenge remains – the ability to identify and track concept drift. While models such as Dish-TS (Fan et al., 2023), FSNet (Pham et al., 2022), and the ensemble-based SOTA method OneNet (Wen et al., 2024) incorporate mechanisms to handle concept drift, they primarily focus on mitigating its impact on forecasting rather than tackling the challenge of adaptive concept identification and evolution.
@@
The evolving study has steered the field towards more adaptive and dynamic models. Methods like change point detection (Deldari et al., 2021; Liu et al., 2023) and online learning algorithms (Zhang et al., 2024) are designed to detect shifts in concepts. However, these methods are typically restricted to detecting structural breaks or focusing on univariate series, rather than tracking and predicting subtle, ongoing changes in concepts, which limits their applicability. In the complex environments, where multiple time series evolve and interact simultaneously, capturing the nonlinear relationships among variables is critical (Marcotte et al., 2023; Bayram et al., 2022). Despite recent advancements (Matsubara & Sakurai, 2019; Li et al., 2022), most models define the concept as a collective behavior of data, falling short in their ability to capture the dynamics of individual series and their interactions. This limitation impairs the models’ capability to discover and interpret the complex interdependencies among variables, thereby constraining their effectiveness in scenarios that involve co-evolving time series, such as city-wide electricity usage or financial market regime detection. This raises a question:
Can we model multiple time series as an evolving ecosystem to identify, track, and forecast concepts, including those unseen in certain series?
We introduce CORAL, a simple yet effective method designed for the dynamic complexities of co-evolving time series data. CORAL employs a kernel-induced self-representation method, adept at capturing the intricate interdependencies and the evolving natures of such data. Our approach, which transforms time series into a matrix format capable of adapting to concept drift, offers a robust and flexible solution for co-evolving time series analysis amidst non-linear interactions and shifting distributions.
@@
(O1) Concept identification: Fig. 1 (b) displays the heatmap of CORAL-generated representation matrix ($n\times n$) over time [^1] CORAL autonomously determines the optimal segments in a domain-agnostic manner.. This matrix showcases a block diagonal structure, where each block represents a unique concept (e.g., C1 - C5), with brightness in the heatmap indicating series correlation strength. Variations of the block structure over time underscore the dynamic correlations within the time series, indicating the presence and evolution of concepts. An extended analysis of these matrices facilitates the identification of the total number of concepts and their pattern (Fig. 1 (c)).
(O2) Dynamic concept drift: Tracking the transitions of time series within the CORAL-generated matrices provides insights into the trajectory of concept drift. Fig. 1 (d) illustrates the concept drift process over time for each of the 3 example series. Red dashed lines in the figure mark the points where concept drift occurs, exemplified by the transition of Series 1 from Concept 4 to Concept 2 and subsequent shifts. This process can also be observed in the representation matrix shown in Fig. 1 (b), where the red star marks the position of Series 1 (S1), and the purple dashed lines trace how S1 shifts over time, further illustrating the process of tracking its transition between concepts.
(O3) Forecasting: The grey areas in Fig. 1 (d) represent the forecasted concept distribution and values of series. These forecasts, denoted by dashed lines matching the colors of the concepts, align closely with the actual series trend (solid grey lines). A significant forecast for Series 1 includes the emergence of Concept 1, previously undetected. This predictive capability stems from CORAL’s ecosystem perspective of the co-evolving time series, leveraging a probabilistic model of the nonlinear interactions among series to anticipate the emergence and evolution of new concepts.
@@
Contributions. CORAL represents more than a mere solution – it is a paradigm shift in co-evolving time series analysis. This framework introduces a new perspective on modeling and interpreting complex, interrelated co-evolving time series. CORAL has the following desirable properties:
1. CORAL introduces a novel, kernel-induced approach for modeling complex interdependencies, enhancing understanding of underlying dynamics, and can be easily integrated into most deep learning backbones.
2. CORAL is adaptive, with the capability to identify and respond to concept drift autonomously, without prior knowledge about concept.
3. CORAL transforms concept drift into a matrix optimization problem, enhancing interpretability and offering a new perspective into the dynamics of co-evolving time series.
@@ 2 The landscape of Concept Drift
Concept Drift. Concept drift in time series refers to the scenario where the statistical properties of the target variable, or the joint distribution of the input-output pairs, change over time. These drifts primarily exhibit in two manners (Ren et al., 2018; Kim et al., 2021): the first is characterized by subtle, ongoing changes, reflecting the evolving dynamics of the time series, while the second arises from sudden shifts caused by structural breaks in the relationships among time series. Understanding dynamic concepts is crucial for complex systems like financial markets, healthcare, and online activity logs, as they reveal underlying structures that improve decision-making and subsequent tasks.
Challenges in Co-evolving Scenarios. Recent advances in time series analysis have led to progress in addressing concept drift. Dish-TS (Fan et al., 2023) alleviates distribution shifts by normalizing model inputs and outputs, while Cogra (Miyaguchi & Kajino, 2019) employs the Sequential Mean Tracker (SMT) to adjust to evolving data distributions. However, these methods struggle with co-evolving time series, where interdependencies cause shifts in one variable to propagate across the system. DDG-DA (Li et al., 2022) adapts data distribution generation for co-evolving scenarios. Yet, it defines concepts as collective behaviors rather than capturing individual series dynamics and interactions. Even the latest deep learning methods, such as OneNet (Wen et al., 2024) and FSNet (Pham et al., 2022), primarily focus on mitigating concept drift’s impact on forecasting rather than tackling adaptive concept identification and dynamic drift tracking. They achieve this through model ensembling or refining network parameters for improved adaptability. Due to the space limit, more related works about concept drift, representation learning on times series and motivation are provided in the Appendix A.
@@ 3 Preliminaries
Concept. Throughout this paper, a concept is defined as the profile pattern of a cluster of similar subseries, observed within a specific segement/window. Here, the term “profile pattern” refers to a subseries, the vector representation of which aligns with the centroid of similar subseries. We use a tunable hyperparameter $\rho$, to differentiate profile patterns and modulate whether concept drift is gradual (smaller $\rho$, more concepts) or abrupt (larger $\rho$, fewer concepts). For a detailed definition, see Appendix A.
Notation. We denote matrices by boldface capital letters, e.g., $\mathbf{M}$. $\mathbf{M}^{\mathrm{T}}$, $\mathbf{M}^{-1}$, Tr$(\mathbf{M})$ indicate transpose, inverse and trace of matrix $\mathbf{M}$, respectively. $\mathrm{diag}(\mathbf{M})$ refers to a vector with its $i$-$th$ element being the $i$-$th$ diagonal element of $\mathbf{M}$.
@@ 3.1 Self-representation Learning in Time Series
Time series often manifest recurring patterns. One feasible way to capture these inherent patterns is through self-representation learning (Bai & Liang, 2020). This approach models each series as a linear combination of others, formulated as $\mathbf{S=SZ}$ or $S_{i}=\sum_{j}S_{j}Z_{ij}$, where $\mathbf{Z}$ is the self-representation coefficient matrix. In multiple time series, high $Z_{ij}$ values indicate similar behaviors or concepts between $S_{i}$ and $S_{j}$. The learning objective function is:
@@ 3.2 Kernel Trick for Modeling Time Series
Addressing nonlinear relationships in co-evolving time series, especially in the presence of concept drift, can be challenging for linear models. Kernelization techniques overcome this by mapping data into higher-dimensional spaces using suitable kernel functions. This facilitates the identification of concepts within these transformed spaces. The process is facilitated by the “kernel trick”, which employs a nonlinear feature mapping, $\Phi(\mathbf{S})$: $\mathcal{R}^{d}\rightarrow\mathcal{H}$, to project data $\mathbf{S}$ into a kernel Hilbert space $\mathcal{H}$. Direct knowledge of the transformation $\Phi$ is not required; instead, a kernel Gram matrix $\boldsymbol{\mathcal{K}}=\Phi(\mathbf{S})^{\top}\Phi(\mathbf{S})$ is used.
4 CORAL
This section introduces the fundamental concepts and design philosophy of CORAL. Our objective is to identify significant concept trends and encapsulate them into a succinct yet powerful and adaptive representative model.
@@ 4.1 Kernel-Induced Representation Learning
$\min_{\mathbf{Z}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{Z}||^{2}$ (2)
$=\min_{\mathbf{Z}}\frac{1}{2}\mathrm{Tr}(\boldsymbol{\mathcal{K}}-\alpha\boldsymbol{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\boldsymbol{\mathcal{K}}\mathbf{Z})$
$s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$
Here, the mapping function $\mathrm{\Phi}(\cdot)$ needs not be explicitly identified and is typically replaced by a kernel $\boldsymbol{\mathcal{K}}$ subject to $\boldsymbol{\mathcal{K}}=\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})$. It’s noteworthy that the parameter $\alpha$ is key to preserving the local manifold structure of time series during this projection, further explained in Sec. 5.2.
@@
$\min_{\mathbf{Z}}\frac{1}{2}$ $\mathrm{Tr}(\boldsymbol{\mathcal{K}}-\alpha\boldsymbol{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\boldsymbol{\mathcal{K}}\mathbf{Z})+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},$ (3)
$s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$
where $\gamma>0$ balances the loss function with regularization term, $||\mathbf{Z}||_{\boxed{\scriptstyle k}}=\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ and $\lambda_{i}(\mathbf{L_{Z}})$ contains the eigenvalues of Laplacian matrix $\mathbf{L_{Z}}$ corresponding to $\mathbf{Z}$ in decreasing order. Here, the regularization term is equal to 0 if and only if $\mathbf{Z}$ is $k$-block diagonal (see Theorem 4.1 for details). Based on the learned high-quality matrix $\mathbf{Z}$ (containing the block diagonal structure), we can easily group the time series into $k$ concepts using traditional spectral clustering technology (Ng et al., 2001). The detailed method of estimating the number of concepts $k$ is provided in Appendix B. To solve Eq. 3, which is a nonconvex optimization problem, we leverage the Augmented Lagrange method with Alternating Direction Minimization strategy to propose a specialized method for solving the nonconvex kernel self-representation optimization (see Appendix C.1).
Theorem 4.1.
@@
Proof.
Due to the fact that $\mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0$, the corresponding Laplacian matrix $\mathbf{L_{Z}}$ is positive semidefinite, i.e., $\mathbf{L_{Z}}\succeq 0$, and thus $\lambda_{i}(\mathbf{L_{Z}})\geq 0$ for all $i$. The optimal solution of $\min\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ is that all elements of $\lambda_{i}(\mathbf{L_{Z}})$ are equal to 0, which means that the $k$ smallest eigenvalues are 0. Combined with the Laplacian matrix property, it’s evident that the multiplicity $k$ of the zero eigenvalues in Laplacian matrix $\mathbf{L_{Z}}$ matches the count of connected components (or blocks) present in $\mathbf{Z}$, and thus the soundness of Theorem 4.1 has been proved. ∎